OpenStudy (anonymous):

For the Future Flvs Students- Algebra 2 Assessment 3.07

OpenStudy (anonymous):

If a baseball player hits a baseball from 4 feet off the ground with an initial velocity of 64 feet per second, how long will it take the baseball to hit the ground? Use the equation h = -16t2 + 64t + 4. (Right Answer) 2 plus or minus square root of 17 end root over 2 quantity of 2 plus or minus square root of 17 all over 2 2 plus or minus 4 square root of 17 quantity of 16 plus or minus square root of 17 all over 2

OpenStudy (anonymous):

Solve 3x2 + 4x = 2. quantity of 2 plus or minus square root of 10 all over 6 quantity of negative 2 plus or minus 2 square root of 10 all over 3 (Right Answer) quantity of negative 2 plus or minus square root of 10 all over 3 quantity of negative 4 plus or minus square root of 10 all over 3

OpenStudy (anonymous):

Which of the following is a solution of x2 + 5x = -2? 5 plus or minus the square root of 33 divided by two. 5 plus or minus the square root of 17 divided by two negative 5 plus or minus the square root of 33 divided by two. (Right Answer) negative 5 plus or minus the square root of 17 divided by two.

OpenStudy (anonymous):

Which of the following is a solution of 4x2 = - 9x - 4? 9 plus or minus the square root of 17 divided by 8. negative 9 plus or minus the square root of 145 divided by 8. 9 plus or minus the square root of 145 divided by 8. (Right Answer) negative 9 plus or minus the square root of 17 divided by 8

OpenStudy (anonymous):

Solve x2 + 6x + 7 = 0. x = -1 and x = -5 3 plus or minus square root of 2 (Right Answer) negative 3 plus or minus square root of 2 quantity of negative 3 plus or minus square root of 2 all over 2